# On point spectrum of Jacobi matrices generated by iterations of quadratic polynomials

### v1

@@ A.3. Computations for $h_{k}(x)$ in a general case
$({\mathfrak{p}}_{2^{n-k}{\varkappa}}(T^{\circ j}(0))\cdots{\mathfrak{p}}_{2^{n-k+j-1}{\varkappa}}(T(0))\Phi_{2^{n-k+j}{\varkappa}})^{*}\Upsilon=\frac{a_{2^{n-k}{\varkappa}-1}\cdots a_{2^{n-k+j}{\varkappa}-1}}{T^{j}(0)\cdots T(0)}\Upsilon$
and
$({\mathfrak{p}}_{2^{n-k}{\varkappa}}(T^{\circ(k+1)}(0))\cdots{\mathfrak{p}}_{2^{n}{\varkappa}}(T(0)))^{*}\Upsilon=\frac{a_{2^{n-k}{\varkappa}-1}\cdots a_{2^{n}{\varkappa}-1}}{T^{k+1}(0)\cdots T(0)}\Upsilon$
we obtain
$f_{0}^{n-(k+1)}=$ $w_{0}^{k+1}(T(0))\frac{a^{2}_{2^{n-k}{\varkappa}-1}\cdots a^{2}_{2^{n}{\varkappa}-1}}{(T^{k+1}(0)\cdots T(0))^{2}}$
$+$ $\sum_{j=0}^{k}w_{0}^{j}(T(0))\frac{a^{2}_{2^{n-k}{\varkappa}-1}\cdots a^{2}_{2^{n-k+j}{\varkappa}-1}}{(T^{j}(0)\cdots T(0))^{2}}f_{j}^{n-k}$
$=$ $w_{0}^{k+1}(T(0))\frac{a^{2}_{2^{n-k}{\varkappa}-1}\cdots a^{2}_{2^{n}{\varkappa}-1}}{(T^{k+1}(0)\cdots T(0))^{2}}$
$+$ $\sum_{j=0}^{k}w_{0}^{j}(T(0))\frac{a^{2}_{2^{n-k}{\varkappa}-1}\cdots a^{2}_{2^{n-k+j}{\varkappa}-1}}{(T^{j}(0)\cdots T(0))^{2}}f_{0}^{n-k+j}.$
With changes of variables we get
$f_{0}^{n-j}=$ $w_{0}^{j}(T(0))\frac{a^{2}_{2^{n-j+1}{\varkappa}-1}\cdots a^{2}_{2^{n}{\varkappa}-1}}{(T^{j}(0)\cdots T(0))^{2}}$
$+$ $\sum_{s=0}^{j-1}w_{0}^{j-s-1}(T(0))\frac{a^{2}_{2^{n-j+1}{\varkappa}-1}\cdots a^{2}_{2^{n-s}{\varkappa}-1}}{(T^{j-s-1}(0)\cdots T(0))^{2}}f_{0}^{n-s}.$ (A.10)
We define vectors
${\mathfrak{f}}^{n}_{k}=\begin{pmatrix}f_{0}^{n}\\ \vdots\\ f^{n-k}_{0}\end{pmatrix},\quad{\mathfrak{g}}^{n}_{k}=\begin{pmatrix}g_{0}^{n}\\ \vdots\\ g^{n-k}_{0}\end{pmatrix},$
where
$g_{0}^{n}=f_{0}^{n}=w_{0}^{0}(T(0)),\quad g_{0}^{n-j}=w_{0}^{j}(T(0))\frac{a^{2}_{2^{n-j+1}{\varkappa}-1}\cdots a^{2}_{2^{n}{\varkappa}-1}}{(T^{j}(0)\cdots T(0))^{2}},\quad 1\leq j\leq k.$
Also we define a lower-triangular matrix ${\mathfrak{L}}^{n}_{k}$ with the entries
$L_{js}=w_{0}^{j-s-1}(T(0))\frac{a^{2}_{2^{n-j+1}{\varkappa}-1}\cdots a^{2}_{2^{n-s}{\varkappa}-1}}{(T^{j-s-1}(0)\cdots T(0))^{2}},\quad 0\leq s\leq j-1.$
Then, due to (A.3), for the vector ${\mathfrak{f}}_{k}^{n}$ we get the system
$(I-{\mathfrak{L}}_{k}^{n}){\mathfrak{f}}^{n}_{k}={\mathfrak{g}}_{k}^{n}.$
It is evident that the system is solvable
${\mathfrak{f}}^{n}_{k}=(I+{\mathfrak{L}}_{k}^{n}+\dots+({\mathfrak{L}}_{k}^{n})^{k}){\mathfrak{g}}_{k}^{n}.$
Moreover we get $f^{n-j}_{0}>0$ for all $j$. As soon as the system is solved we have
${\mathcal{M}}_{2^{n-k}{\varkappa}}\cdots{\mathcal{M}}_{2^{n-1}{\varkappa}}{\mathcal{M}}_{2^{n}{\varkappa}}w_{x}^{0}(T(0))I=w_{x}^{0}(T(0))f^{n-k}_{0}\Phi_{2^{n-k}{\varkappa}}\star I$
$+w_{x}^{1}(T(0))f^{n-k+1}_{0}({\mathfrak{p}}_{2^{n-k}{\varkappa}}(T(0))\Phi_{2^{n-k+1}{\varkappa}})\star I+\dots$
$+w_{x}^{j}(T(0))f^{n-k+j}_{0}({\mathfrak{p}}_{2^{n-k}{\varkappa}}(T^{\circ j}(0))\cdots{\mathfrak{p}}_{2^{n-k+j-1}{\varkappa}}(T(0))\Phi_{2^{n-k+j}{\varkappa}})\star I$
$+\dots+w_{x}^{k}(T(0))f^{n}_{0}({\mathfrak{p}}_{2^{n-k}{\varkappa}}(T^{\circ k}(0))\cdots{\mathfrak{p}}_{2^{n-1}{\varkappa}}(T(0))\Phi_{2^{n}{\varkappa}})\star I$
$+w_{x}^{k+1}(T(0))({\mathfrak{p}}_{2^{n-k}{\varkappa}}(T^{\circ(k+1)}(0))\cdots{\mathfrak{p}}_{2^{n-1}{\varkappa}}(T^{\circ 2}(0)){\mathfrak{p}}_{2^{n}{\varkappa}}(T(0)))\star I$

### v2

@@ 4.2. Equivalence Lemma
${\mathfrak{L}}_{n+1,i}({\varkappa})=w^{n-i}_{0}(T(0))\langle{\mathfrak{P}}^{n-i}_{\hat{}\underline{s}^{i}{\varkappa}}(T(0))\Psi_{\hat{}\underline{s}^{i-1}{\varkappa}},\Upsilon_{\hat{}\underline{s}^{n}{\varkappa}}\rangle^{2}$
the right hand sides in (4.7) and (4.8) coincide and we get the following interpolation property for $f^{n}_{0}({\varkappa})$
$f^{n+1}_{0}({\varkappa})=\langle h_{n}(0)\Upsilon_{\hat{\underline{s}}^{n}{\varkappa}},\Upsilon_{\hat{\underline{s}}^{n}{\varkappa}}\rangle.$ (4.9)
Consequently $f^{n}_{0}({\varkappa})\geq C_{1}$ and the lemma is proved. ∎
4.3. Uniform boundedness from below
First let us mention compactness arguments. Let $K$ be a compact (symmetric ellipse) containing the interval $\xi,\xi]$ which does not contain $\lambda$. Then for all $n\geq n_{0}(K)$ and $z\in K$
$\frac{1}{2}\leq\left|\frac{z-T^{\circ n}(0)}{T^{\circ n}(0)}\right|\leq 2.$
Due to the relation
$w_{z}^{n}(T(0))=\frac{-T^{\circ(n+1)}(0)}{z-T^{\circ(n+1)}(0)}w_{0}^{n}(T(0)),$
we can extend the uniform estimation of $w_{z}^{n}(T(0))$ for all $z\in K$ using (3.8). By (3.7)
$\|h_{n}(z)\|\leq C(K)\sum_{k=0}^{n}\text{\rm tr}\,A^{n}_{k},$
and by (3.10) $\|h_{n}(z)\|\leq C_{1}(K)$.
Now we mention a kind of weak convergence for the sequence $\{f_{0}^{n}\}$ in the class ${\mathcal{F}}_{N}$.
Lemma 4.3.
Let ${\varkappa}\in{\mathcal{F}}_{N}$. Assume that $f^{n}_{0}\to 0$ along a certain subsequence of indices $\tilde{I}$. then for an arbitrary fixed $n_{0}$ and a sequence $k(n)\in{\mathbb{N}}$ with $k(n)\leq n_{0}$ we have
$\lim\limits_{\begin{subarray}{c}n\in\tilde{I}\end{subarray}}f^{n-k(n)}_{0}=0.$
Proof.
This follows from the estimate
$f^{n}_{0}\geq{\mathcal{L}}_{n,n-1}f_{0}^{n-1}=w_{0}^{0}(T(0))\langle\Psi_{\hat{\underline{s}}^{n-2}{\varkappa}},\Upsilon_{\hat{\underline{s}}^{n-2}{\varkappa}}\rangle f_{0}^{n-1}.$
Therefore, by (4.4), $f^{n}_{0}\geq Cf_{0}^{n-1}$ with a constant $C>0$ depending only on $N$. ∎
Lemma 4.4.
Let ${\varkappa}\in{\mathcal{F}}_{N}$. Then there exists $C>0$ such that $h_{n}(x)\geq C\cdot I$ for all $n$ and $x\in[-\xi,\xi]$.
Proof.
Assuming the contrary, due to Lemma 4.2 (specifically see (4.9)), we have a subsequence of indices $\tilde{I}$ such that
$\lim_{\begin{subarray}{c}n\in\tilde{I}\end{subarray}}f^{n+1}_{0}=\lim_{n\in\tilde{I}}\langle h_{n}(0)\Upsilon_{\hat{\underline{s}}^{n}{\varkappa}},\Upsilon_{\hat{\underline{s}}^{n}{\varkappa}}\rangle=0.$ (4.10)
By Lemma 2.6 and Lemma 4.3 we can pass to another subsequence $I$ such that $\hat{\underline{s}}^{n}{\varkappa}$ is odd, $\hat{\underline{s}}^{n-1}{\varkappa}$ is even and (4.10) still holds for $n\in I$.
Due to compactness of ${\mathbb{Z}}_{2}$ we have $\varkappa$$^{\prime}$ and a subsequence $I_{{\varkappa}^{\prime}}\subset I$ such that
$\lim_{n\in I_{{\varkappa}^{\prime}}}\hat{\underline{s}}^{n}{\varkappa}={\varkappa}^{\prime}.$ (4.11)
Finally, due to compactness of $H^{\infty}(K)$ we choose a subsequence $I^{\prime}_{{\varkappa}^{\prime}}\subset I_{{\varkappa}^{\prime}}$ such that for some $h_{\infty}\in H^{\infty}(K)$, uniformly on $K$,
$\lim_{n\in I^{\prime}_{{\varkappa}^{\prime}}}h_{n}(z)=h_{\infty}(z).$ (4.12)
We will show that the matrix function $h_{\infty}(z)$ is of the form
$h_{\infty}(z)=\begin{pmatrix}h^{\infty}_{11}(z)&0\\ 0&0\end{pmatrix}$. (4.13)
We note that the function $w_{x}^{0}(T(0))=1/(x+\lambda)$ is decreasing on $x\in[-\xi,\xi]$, but that for $k\geq 1$, the function $w_{x}^{k}(T(0))$ is increasing on $x\in[-\xi,\xi]$, because $T^{\circ j}(0)>0$ for $j\geq 2$. For this reason, in (3.7) we separate the $k=0$ term and introduce
$h^{\sharp}_{n}(x)=\sum_{k=1}^{n}w^{k}_{x}(T(0))A^{n}_{k}.$
By Lemma 4.3, (4.10) implies $f^{n}_{0}\to 0$ for $n\in I^{\prime}_{{\varkappa}^{\prime}},n\to\infty$. Since by (4.1),
$A^{n}_{0}=f^{n}_{0}({\varkappa})\Psi_{\hat{}\underline{s}^{n-1}{\varkappa}}\Psi_{\hat{}\underline{s}^{n-1}{\varkappa}}^{*}$
we get by (4.12)
$\lim_{n\in I^{\prime}_{{\varkappa}^{\prime}}}h^{\sharp}_{n}(z)=h_{\infty}(z).$
Since $h^{\sharp}_{n}(x)$ are positive and increasing (in $x$) on $\xi,\xi]$ we get that $h_{\infty}(x)$ is non-negative and non-decreasing in this interval. By our choice of $I$ all terms $\hat{\underline{s}}^{n}{\varkappa}$ are odd. Therefore
$\Upsilon_{\hat{\underline{s}}^{n}{\varkappa}}=\Upsilon_{{\varkappa}^{\prime}}=\begin{pmatrix}0\\ 1\end{pmatrix}$.
By (4.10), we have
$\langle h_{\infty}(0)\Upsilon_{{\varkappa}^{\prime}},\Upsilon_{{\varkappa}^{\prime}}\rangle=0.$
By non-negativity and non-decreasing property we get
$\langle h_{\infty}(x)\Upsilon_{{\varkappa}^{\prime}},\Upsilon_{{\varkappa}^{\prime}}\rangle=0,$ (4.14)
for all $x\in[-\xi,0]$. By analyticity (4.13) holds for all $x\in[-\xi,\xi]$. Non-negativity of $h_{\infty}(x)$ implies that off-diagonal entries also vanish. Thus (4.13) is proved.
Now we use that $\hat{\underline{s}}^{n-1}{\varkappa}$ is even. Note that in particular
$\lim_{n\in I^{\prime}_{{\varkappa}^{\prime}}}\hat{\underline{s}}^{n-1}{\varkappa}=2{\varkappa}^{\prime}.$
By passing to a further subsequence $I^{\prime\prime}_{{\varkappa}^{\prime}}\subset I^{\prime}_{{\varkappa}^{\prime}}$ so that
$\tilde{h}_{\infty}(x)=\lim_{n\in I^{\prime\prime}_{{\varkappa}^{\prime}}}h_{n-1}(x),$
and repeating the above arguments, we use
$\lim_{n\in I^{\prime\prime}_{{\varkappa}^{\prime}}}\langle h_{n-1}(0)\Upsilon_{\hat{\underline{s}}^{(n-1)}{\varkappa}},\Upsilon_{\hat{\underline{s}}^{(n-1)}{\varkappa}}\rangle=\lim_{n\in I^{\prime\prime}_{{\varkappa}^{\prime}}}f_{0}^{n}=0$
and using $\Upsilon_{\hat{\underline{s}}^{(n-1)}{\varkappa}}=\Upsilon_{2{\varkappa}^{\prime}}=\binom{1}{0}$, we conclude as before that
$\tilde{h}_{\infty}(x)=\begin{pmatrix}0&0\\ 0&\tilde{h}^{\infty}_{22}(x)\end{pmatrix}$.
On the other hand
$h_{n}(x)=({\mathcal{M}}_{\hat{\underline{s}}^{n-1}{\varkappa}}h_{n-1})(x).$
Passing to the limit we have
$h_{\infty}(x)=({\mathcal{M}}_{2{\varkappa}^{\prime}}\tilde{h}_{\infty})(x)=\frac{1}{2}\sum_{T(y)=x}{\mathfrak{p}}_{2{\varkappa}^{\prime}}(y)\begin{pmatrix}0\\ 1\end{pmatrix}\tilde{h}^{\infty}_{22}(y)\begin{pmatrix}0&1\end{pmatrix}{\mathfrak{p}}^{*}_{2{\varkappa}^{\prime}}(y)$
$=\frac{1}{2}\sum_{T(y)=x}\begin{pmatrix}0\\ 1\end{pmatrix}\tilde{h}^{\infty}_{22}(y)\begin{pmatrix}0&1\end{pmatrix}$.
Thus $h^{\infty}_{11}(x)=0$, i.e., $h^{\infty}(x)=0$, which contradicts the uniform bound $\text{\rm tr}\,h_{\infty}(x)\geq C>0$, see (3.13). ∎
4.4. Proof of the Main Theorem
Theorem 4.5.
If ${\varkappa}\in{\mathcal{F}}_{N}$, then the spectrum of $J_{\varkappa}$ is singular continuous.
Proof.
We have
$\int g(T(x))\text{\rm tr}\left(h_{n}(x)\Sigma_{\hat{\underline{s}}^{n}{\varkappa}}(dx)\right)=\int\text{\rm tr}\left(g(T(x))h_{n}(x)({\mathcal{M}}^{*}_{\hat{\underline{s}}^{n}{\varkappa}}\Sigma_{\hat{\underline{s}}^{n+1}{\varkappa}})(dx)\right)$
$\int\text{\rm tr}\left(({\mathcal{M}}_{\hat{\underline{s}}^{n}{\varkappa}}g(T(x))h_{n}(x))\Sigma_{\hat{\underline{s}}^{n+1}{\varkappa}})(dx)\right)=\int g(x)\text{\rm tr}\left(h_{n+1}(x)\Sigma_{\hat{\underline{s}}^{n+1}{\varkappa}})(dx)\right)$
[…]
$\text{\rm tr}\left(h_{n+1}(T^{\circ(n+1)}(x_{0}))\Sigma_{\hat{\underline{s}}^{n+1}{\varkappa}}(\{T^{\circ(n+1)}(x_{0})\})\right)=$
$\text{\rm tr}\left(h_{n}(T^{\circ n}(x_{0}))\Sigma_{\hat{\underline{s}}^{n}{\varkappa}}(\{T^{\circ n}(x_{0})\})\right)+\text{\rm tr}\left(h_{n}(-T^{\circ n}(x_{0}))\Sigma_{\hat{\underline{s}}^{n}{\varkappa}}(\{-T^{\circ n}(x_{0})\})\right)$
On the other hand since
$c_{-}\cdot I\leq h_{n}(x)\leq c_{+}\cdot I,\quad x\in[-\xi,\xi],$
we have
$\text{\rm tr}\left(h_{n}(-y_{0})\Sigma_{\hat{\underline{s}}^{n}{\varkappa}}(\{-y_{0}\})\right)\geq c_{-}\text{\rm tr}\Sigma_{\hat{\underline{s}}^{n}{\varkappa}}(\{-y_{0}\})=$
$c_{-}\text{\rm tr}\Sigma_{\hat{\underline{s}}^{n}{\varkappa}}(\{y_{0}\})\geq\frac{c_{-}}{c_{+}}\text{\rm tr}\left(h_{n}(y_{0})\Sigma_{\hat{\underline{s}}^{n}{\varkappa}}(\{y_{0}\})\right).$
(in this step we used that the trace of the spectral matrix measure is even, since all diagonal Jacobi coefficients are zero). In combination we get
$\text{\rm tr}\left(h_{n+1}(T^{\circ(n+1)}(x_{0}))\Sigma_{\hat{\underline{s}}^{n+1}{\varkappa}}(\{T^{\circ(n+1)}(x_{0})\})\right)\geq$
$(1+\delta)\text{\rm tr}\left(h_{n}(T^{\circ n}(x_{0}))\Sigma_{\hat{\underline{s}}^{n}{\varkappa}}(\{T^{\circ n}(x_{0})\})\right)$
with $\delta=c_{-}/c_{+}$. If $\text{\rm tr}\,\Sigma_{\varkappa}(\{x_{0}\})>0$, then $\text{\rm tr}\,\left(h_{0}(x_{0})\Sigma_{\varkappa}(\{x_{0}\})\right)>0$ and the sequence $\text{\rm tr}\left(h_{n}(T^{\circ n}(x_{0}))\Sigma_{\hat{\underline{s}}^{n}{\varkappa}}(\{T^{\circ n}(x_{0})\})\right)$ is unbounded. On the other hand, the measures $\Sigma_{\hat{\underline{s}}^{n}{\varkappa}}$ and functions $h_{n}(x)$ are uniformly bounded. ∎
Proof of Theorem 1.1.
Since ${\mathbb{Z}}_{2}={\mathcal{G}}_{\delta}\cup{\mathcal{F}}_{\sigma}$ and ${\mathcal{F}}_{\sigma}=\cup_{N}{\mathcal{F}}_{N}$, this follows from Theorems 3.5 and 4.5. ∎